3. c) A wire of length 7 m and radius 2.37 mm is stretched by a load of 12370 gm and increased by 1.78% of the original length. Find the stress, strain, and Young's modulus of the material of the wire.

1
Expert's answer
2021-12-06T16:42:59-0500
"\epsilon=0.0178\\\sigma=\frac{12.37\cdot9.8}{\pi\cdot(2.37)^2\cdot10^{-6}}=6.87\cdot10^{6}\ Pa\\Y=\frac{6.87\cdot10^{6}}{0.0178}=3.86\cdot10^{8}Pa"